### v2

# A 44-minute periodic radio transient in a supernova remnant

Abstract
[…]
Keyword: radio astronomy; long period radio transient; supernova remnant; radio pulsar
These authors contributed equally to this work
* Corresponding authors: yanjingye@nssc.ac.cn
I Introduction
Long-period radio transients (LPTs) are a new class of periodic radio emitters, characterized by ultra-long rotation periods and strong magnetic fields. These features place LPTs significantly outside the conventional framework of rotation-powered neutron star pulsars or white dwarfs (marsh2016radio; pelisoli20235; hurley2022radio; hurley2023long; caleb2024emission; rea2024long; de2025sporadic; lee2025emission; dong2025chime; bloot2025strongly). Among reported LPT cases, several are identified, or argued to originate from, white-dwarf binary systems marsh2016radio; pelisoli20235; de2025sporadic; hurley20242; dong2025chime; bloot2025strongly, including both spin-down powered white-dwarf pulsars (i.e. AR Sco) and accretion-powered systems. Another group shows no compelling evidence for companions: GLEAM-X J162759.5-523504.3 (18.18 min) hurley2022radio, GPM J1839-10 (21 min) hurley2023long, ASKAP J1935+2148 (53 min) caleb2024emission, CHIME J0630+25 (7 min) dong2024discovery. For the stellar candidates LPTs without confirmed companions, the two primary models—spin-down powered pulsars and white dwarfs—each face significant challenges in explaining the observed emissions if they independently evolve in a traditional way: losing rotation energy through radio emission. Neutron stars with ultra-long periods typically fall below the so-called “death line”, where their rotational energy becomes insufficient to sustain coherent radio emission chen1993pulsar; young1999radio; caleb2022discovery, while white dwarfs often have insufficient magnetic fields to produce the observed periodic, coherent emissions rea2024long. Although theoretical models lean toward isolated LPTs being magnetars (i.e. neutron stars with strong magnetic field) cooper2024beyond; men2025highly; beniamini2023evidence, direct observational evidence and evolution studies remain limited.
II Observation and Results
II.1 DART Observation
We report a new LPT in a supernova remnant, DART J1832-0911, detected by DAocheng Radio Telescope (DART) (see supplementary material) with interferometric imaging across a frequency range of 149-459 MHz. The first pulse was detected on Modified Julian Date (MJD) 60374. From MJD 60401 to MJD 60443, we conducted continuous observations of the transient using an interferometric imaging technique with full polarization, simultaneously recording data at two distinct center frequencies, each with a bandwidth of 4 MHz. The recorded pulses showed an estimated peak flux density between 0.5–2 Jy. After this period, the transient entered a long-period quiescent state, as recorded by DART observation.
The interferometric image allowed us to determine the celestial coordinates to be RA=18h32m48.5s () and DEC=-09d11m17s (), situating it within the projected region of supernova remnant G22.7-0.2. Top panel in Fig.1 displays the pulse-on image of DART J1832-0911 at 425 MHz. The bottom panel shows the observed single pulses in two Stokes components. Additional pulse profile shapes are provided in the supplementary material (Fig.S1). All pulse series presented in this study have been aligned through periodic folding and corrected for delay time. By analyzing the time of arrival (TOA) of single pulses recorded by DART, we derive a spin period of s with an estimated period derivative of s/s. Fig.2 displays the - diagram of DART J1832-0911 as well as other LPTs.
II.2 FAST Observation
To further investigate, we used the Five-hundred-meter Aperture Spherical Telescope (FAST) to observe the source at L-band during DART’s non-detection period on MJD 60559. During a 1-hour observation, we detected a short pulse with a peak flux density of approximately 94 mJy. The duration of the pulse (i.e. the pulse width) is 0.2 s, just 0.007% of the rotation period. This could point to a highly energetic and sharp emission event occurring at the strong magnetic field, like the dwarf pulse emission process of a neutron star chen2023strong; yan2024dwarf.
We folded the time series based on the ephemeris derived from DART observations and found that the TOA of this pulse is within the predicted emission phase. This short pulse is shown in Fig.3. The rotation measure (RM) of this pulse is . A summary of the measured parameters of this transient is provided in Table.1.
| Parameter | DART | FAST |
| 149-459 MHz | 1-1.5 GHz | |
| Period derivative (s/s) | - | |
| DM () | 480 | 464.5 |
| RM () | - | |
| Pulse Width (s) | 40-250 | 0.2 |
| Peak flux density (Jy) | 0.5-2 | 0.0960.027 |
II.3 Optical and X-ray Counterpart Search
We conducted an optical observation using the Gran Telescopio Canarias (GTC) on 30 July 2024. The integration time for this observation was 1.25 hours.
No credible optical counterpart was identified in the GTC observation, effectively ruling out any pulsating white dwarf with a temperature 17,000K as its origin. An r-band image of the DART source location was obtained using the broad-band imaging mode of OSIRIS on the GTC. As shown in Fig.4, there are four optical candidates within the designated zone. However, the light curves of these candidates show no periodic variations on timescales of tens of minutes, indicating they are not associated with DART J1832-0911. Given an upper limit of 26.6 magnitudes for the GTC observations and assuming a white dwarf radius of at a distance of 4.5 kpc, any WD detectable by the GTC would have an effective temperature below approximately , indicating an evolution age of hundreds of millions years (see supplementary material). The absence of an optical counterpart supports the possibility of a compact neutron star rather than a WD. Additionally, our analysis shows that a spin-down powered WD would be unable to account for the observed emission power of a transient with characteristics like those of DART J1832-0911 (see supplementary material).
We searched for any possible X-ray counterpart using archival data from the XMM-Newton observatory, obtained on March 13, 2011 (Observation ID: 0654480101). The observation had an exposure time of s and covered the source with both XMM-PN and XMM-MOS2 detectors in the 0.2-10 keV. Within the positional uncertainty of DART J1832-0911, only a single photon was detected by XMM-PN. Consequently, no X-ray counterpart was identified for either the transient or the associated SNR region according to the observation in 2011. This yields an upper limit of erg cm-2 s-1 in 0.2-10 keV at a 99% confidence level.
Independently of our work, a recent study wang2025detection has reported a 44-min periodic radio transient in the same direction, consistent with our detection (referred to as ASKAP J18320911 in their work). The authors performed a comprehensive X-ray analysis of J18320911 using additional instruments, including Chandra, Swift, and Einstein Probe Yuan2015EP, leading to the identification of a transient X-ray counterpart detected by Chandra. Their observations revealed an extreme contrast between a bright X-ray outburst ( erg s-1) and a deep quiescent state ( erg s-1). Assuming a distance of 4.5 kpc, these correspond to flux limits of approximately erg cm-2 s-1 in the 1–10 keV band. We suggest that rapid follow-up X-ray observations conducted during the radio-bright phases of future detected LPTs will be critical for unveiling their physical nature.
Archival XMM-Newton observations from 2011 revealed no detectable X-ray emission from DART J1832–0911, whereas an X-ray burst was detected by Chandra in February 2024 but not in August (wang2025detection), indicating that J1832–0911 is an intermittent X-ray transient.
II.4 Dispersion Measure and SNR Association Analysis
We calculated the dispersion measure (DM) using dual-frequency pulse series from DART and the short pulse spectra of FAST. By relating the TOA delay to the dispersion effect observed at two frequencies simultaneously recorded, we obtained a DM of pc . A DM search of the FAST spectra yielded . These measurements from the two observatories are consistent within the 3 sigma error. Given the broader bandwidth and finer temporal resolution of FAST, we consider this result to be the precise DM for DART J1832-0911.
We estimated the DM distance of DART J1832-0911, it is consistent with the distance of SNR G22.7-0.2. Previous studies estimate the distance to SNR G22.7-0.2 at 4.4 kpc su2014interaction or 4.7 kpcranasinghe2018revised, based on molecular cloud analyses. Meanwhile, the DM-derived distance to DART J1832-0911 was estimated using three Galactic free-electron density models: YMW16 yao2017new, NE2001 cordes2002ne2001, and NE2025 ocker2026ne2025, yielding distances of approximately 4.5 kpc, 5.9 kpc, and 6.4 kpc, respectively. All three distances are broadly consistent with the SNR distance given the known systematic uncertainties of free-electron density models, which can reach 30–50%. This suggests that J1832-0911 may reside within the SNR bubble. Nevertheless, the DM-estimated distance remains debated due to the electron-density model dependence.
To evaluate the reliability of DM-based distance estimation along this line of sight, we compare DM-derived distances with VLBI parallax distances with two pulsars deller2019microarcsecond,which are located within in declination of DART J18320911: PSR J18200427 and PSR J19010906. For PSR J18200427 (DM pc cm-3), the NE2001, YMW16, and NE2025 models yield distances of 1.94, 2.92, and 2.01 kpc, respectively, while a VLBI parallax distance is kpc. For PSR J19010906 (DM pc cm-3), the DM models give 2.13, 2.89, and 2.43 kpc, and the VLBI distance is kpc. We estimate the uncertainty in DM-derived distances based on the VLBI measurements using two approaches. First, we evaluate each of the three DM models individually. The average fractional deviations between the DM distances and the VLBI distances for the two sightlines are 20% for NE2001, 25% for YMW16, and 27% for NE2025. Based on these model-specific uncertainties, the DM-derived distances of DART J18320911 are 4.7–7.1 kpc (NE2001), 3.4–5.6 kpc (YMW16), and 4.7–8.1 kpc (NE2025). Taking the full range of these estimates, we adopt a DM distance of 3.4 to 8.1 kpc. The second estimation method combines all three DM models. We take the mean of the three model predictions as the reference distance, kpc. The total uncertainty is estimated by combining in quadrature the mean single-model systematic uncertainty (24%) and the fractional model-to-model dispersion (), yielding a combined fractional uncertainty of , or 2.4 kpc. This gives a consolidated distance range for DART J18320911 of 3.2–8.0 kpc. The independently determined distance to SNR G22.70.2 of 4.4–4.7 kpc su2014interaction; ranasinghe2018revised falls within both the individual-model ranges and the consolidated interval, supporting a physical association between DART J18320911 and the SNR.
[…]
We inspected the CO observations toward SNR G22.7-0.2. The velocity of CO emission ranges from 75 to 79 km/s su2014interaction, consistent with the absorption ( km/s) reported in Wang et al. wang2025detection. Such alignment in the gas dynamics strongly supports that the compact object and the SNR reside within the same large-scale ISM structure.
We inspect all compact objects within the 16 deg2 region FoV of our DART observation. There are 17 other sources, which fall within the DM distance of 2 kpc to 8 kpc and an age below 1 million years, accommodating a 100% uncertainty allowed for the DM uncertainty and the SNR distance, and a 1000% uncertainty allowed for the upper limits of the SNR age, as indicated in the Fig.5. The random chance of any of these sources to get within the distance between the LPT and the SNR center is lower than 0.7% (see supplementary material). Note that these sources have a chance of being SNR related. The contamination from the field pulsar population has to be much lower. And the previous pulsar (PSR) -SNR association studies reported no case of pulsars discovered around SNR to be later revealed to a chance coincidence (see supplementary material for a detail discussion).
We also investigated the X-ray emission of SNR G22.7-0.2 using data from XMM-Newton and previous studies. Our analysis, consistent with earlier research, reveals no evidence of extended X-ray emission within the radio range . The cessation of X-ray emission from the SNR suggests that the remnant’s shock velocity, , has decreased below 200 km/s—insufficient to heat the gas to X-ray-emitting temperatures mckee1974x; chevalier1999supernova. Consequently, we can reasonably infer that the SNR has entered the radiative phase, during which the age of the SNR can be estimated as years where pc is the SNR radius. Assuming a lower limit for the shock velocity of 30 km/s which is the minimum shock velocity required to produce the observed high-velocity wings and redshifted broadening in the CO emission of this SNR su2014interaction, we obtain an upper age limit of years. Assuming that DART J18320911 is associated with SNR G22.70.2, we use the age and radius of the SNR to estimate its projected transverse velocity to be 26–176 km/s (see Section S6 for details), which aligns well with the velocity range of the SNR-associated pulsarsmalov2021pulsars.
Conclusively, the widely used association judgement: positional overlap, distance, proper motion, and the consistency between the SNR age and the spin-down timescale expected from fallback disk models, shows no evidence against the physical association between DART J1832-0911 and SNR G22.7-0.2. Nevertheless, further strong evidence from direct observational measurements is required .
II.5 Radio Emission Properties
DART J1832-0911 displays a range of emission characteristics during its active radio period. Prior to either undetectable weak emission or nulling, the source undergoes mode changes modulated by variations in pulse width and strength, as illustrated in the bottom panel of Fig.1. These time-dependent mode changes reveal an evolution from bright, wide pulses to weaker, narrower ones. In the wide mode, pulse widths range from 100 s to 300 s, with a flux density of about 2 Jy. In the narrow mode, pulse width narrows to approximately 40 s, and flux density decreases to roughly 0.5 Jy. During the emission-quiescent state, as observed by DART, we captured a short-lived pulse by FAST. Within the expected periodic phase, this pulse lasted 0.2 s and exhibited a peak flux of mJy. Such brief, isolated emission components have also been observed in other long-period transients hurley2023long; caleb2024emission.
The polarization recorded of DART J1832-0911 presents unique properties as it significantly diverges from the other LPTs. While all reported isolated LPTs—except CHIME J0630+25 for which the research provided no polarization information—exhibit notable polarized components throughout their emission cycles, DART’s pulses show a sporadic distribution of polarization components. Additionally, many pulses are extremely poorly polarized throughout the entire phase. Aside from a few sharp impulsive linear polarization (LP) components, we found no other LP emissions with a signal-to-noise ratio (S/N) greater than for most pulses. Fig.S7 highlights two instances of narrow LP components with S/N of 14.2 (MJD 60416) and 6.6 (MJD 60421), respectively. Regarding circular polarization (CP), we observed that pulses in the wide mode exhibit narrow and phase-locked CP components, as illustrated in the bottom panel of Fig.1. The peak flux ratio of the highest CP components to the aligned total intensity reaches V/I. Additionally, most CP components display sign reversal (see Fig.S7). The narrow width of the CP components, spanning less than 3% of the period, emphasizes a fixed circularly polarized emission zone in a limited region. Notably, the short-lived pulse detected by FAST is nearly 100% linearly polarized, and the PA is flat across the whole emission duration. Fig.3 shows the dynamic spectra and polarization properties of this pulse.
III Conclusion
We report the discovery of a long-period radio transient, DART J18320911, detected with interferometric imaging by the Daocheng Radio Telescope (DART) across 149–459 MHz. The spin period of this LPT is s (44.27 min). The interferometric localization (RA , Dec , ) situates DART J18320911 within the projected extent of the supernova remnant G22.70.2. During the DART non-detection epoch, a short pulse detected by FAST at L-band occurred at the predicted rotational phase and is nearly linearly polarized with a well-measured RM, demonstrating that the source can emit extremely narrow, highly polarized bursts despite long intervals of radio quiescence.
The DM measurements of J1832-0911 from DART and FAST are roughly consistent, and the DM-derived distance (, and kpc from YMW16, NE2001, and NE2025 models, respectively) are broadly consistent with independent distance estimates for SNR G22.70.2, supporting a physical association between the compact object and the remnant environment. Complementary CO kinematics and HI absorption arguments further strengthen the case that both the SNR and the transient reside in the same large-scale ISM structure. A field-population analysis yields a low probability for a chance alignment with unrelated compact objects, adding statistical weight to the association scenario.
DART J18320911 shows distinctive radio emission and polarization phenomenology. In DART bands, the activity evolves between wider and narrower emission modes and ultimately into a long quiescent state, with peak flux densities of order 0.5–2 Jy during active epochs. Polarization properties are notably intermittent: many pulses are weakly polarized, while (i) narrow, phase-locked circular polarization components appear in the wide mode with reaching and occasional sign reversals, indicating a stable and geometrically confined emission region; and (ii) the FAST burst shows nearly pure linear polarization with a flat position angle across the burst window, pointing to an ordered magneto-ionic geometry during that event.
Multi-wavelength follow-up and archival searches impose additional constraints. Optical observation reveals no convincing counterpart down to a deep limiting magnitude, disfavoring a luminous white-dwarf binary interpretation. Archival XMM–Newton observations from 2011 yielded no detectable X-ray counterpart, setting a 99% upper limit on the 0.2–10 keV flux. Together with the reported Chandra detection of X-ray bursts and subsequent non-detection, this suggests that the source can enter a deep X-ray quiescent state and motivates rapid, triggered X-ray follow-up during future radio-bright windows.
IV DISCUSSION
DART J1832-0911 provides several important advances over previously reported LPTs. First, it is the strongest current case for a physical association between an LPT and a supernova remnant, thereby providing the clearest observational support so far that at least some isolated LPTs originate from young neutron stars. Second, the plausible SNR association supplies an external age constraint, allowing the ultra-long spin period to be discussed in an evolutionary context and making post-supernova braking scenarios, such as fallback-disk interaction, substantially more plausible. Third, the source exhibits distinctive polarization behaviour, including phase-locked circular polarization and a high linearly polarized dwarf pulse, which places new constraints on the radio-emission geometry and further distinguishes this source from LPTs associated with white-dwarf binary or accretion-powered systems. Taken together, these properties favor a young neutron star as one origin of at least a subset of long-period radio transients.
The spatial association between this source and the SNR indicates that it is likely the stellar remains of a supernova explosion, specifically a neutron star rather than a white dwarf. Given the characteristic age of both DART J1832-0911 and SNR G22.7-0.2, it is reasonable to suggest that DART J1832-0911 might be a young neutron star. The emission properties further suggest a pulsar-like emitter of DART J1832-0911. The narrow, phase-locked CP indicates a stable geometric structure of the magnetic field of DART J1832-0911 even at small scales. This stability argues against models that rely on dynamic magnetic fields as the primary driver of radio emission, such as star flares benz2010physical and dwarf binary systems marsh2016radio; de2025sporadic. The nearly 100% linearly polarized burst signals support both theoretical radio emission mechanism and observation related to magnetarscamilo2006transient; camilo2007polarized; kramer2007polarized; levin2012radio, which is believed to be a class of young neutron stars kaspi2017magnetars. Such short-lived pulse emitted during an emission-dull or -quiescent state is similar to the dwarf pulse observed during pulsars’ nulling periodchen2023strong; yan2024dwarf; srostlik2005core; kloumann2010long; esamdin2012psr. These emissions are believed to arise from rain-like drops of particles formed via pair production near the neutron star surfacechen2023strong; yan2024dwarf. It is worth noting, however, that in the case of LPTs, such emissions may occur regularly during both emission-loud and -quiet phases, but remain undetected due to the sensitivity and temporal resolution limitations of DART.
A neutron star assumption of LPTs significantly challenges conventional evolution models of fast spinning neutron star scenariosrea2024long, which typically have rotation periods of less than a few minutes, as illustrated in Fig.2. Theoretically, neutron stars—formed as remnants of supernova explosions—are born with fast spin periods on the order of milliseconds if they inherit most or all of the angular momentum from their progenitors. Consequently, the emergence of LPTs suggests that these stars could form under two possible conditions: either with a significant loss of angular momentum from their progenitors such as neutrino emission epstein1978neutrino; camelio2016spin, or through some rotational braking process during their spin evolution (see heger2005presupernova for a review).
Given that DART J18320911 is embedded in the SNR environment, rotational braking by a supernova fallback disk menou2001stability; ronchi2022long; yang2024instability provides the most natural explanation for its ultra-long spin period. We simulate the reasonable period range based on a fallback disc model yang2024instability, as shown in Fig.2. The results suggest that the observed LPTs could be well explained in the period between and years after the supernova explosion. Therefore, the association between DART J1832-0911 and SNR G22.7-0.2, with an age estimate ranging from to years, could further support the hypothesis that such LPTs may originate from interactions between young neutron stars and fallback discs.
Acknowledgements.
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences.Supplementary data for
A 44-minute periodic radio transient in a supernova remnant
S1 DART Observation & Data Reduction
[…]
For the majority of the data, we configure the integration time within a range from 100 to 500 milliseconds. For each observation, two center frequencies are selected from within the entire frequency band. Ultimately, a pulse series spanning from 190 MHz to 455 MHz is recorded. All these pulses are acquired with complete Stokes components. Fig.S1 showcases a portion of the pulse profile components.
In DART system, the polarimetric measurement accuracy of DART is achieved by amplitude and phase unbalance calibration with the external and internal calibration. The calibration refers to the correction of instrumental gain and phase mismatches between all 626 channels. Specifically, the measurement is composed by direct dual-pol observation(HH,VV) and full-polarimetric cross-correlation (HH,VV and HV). A post-data processing can finally convert the polarimetric measurement to the Stokes parameters as the form like
We illustrate the data processing chain in Fig.S2. The polarimetric measurement accuracy only depends on the amplitude and phase unbalance of the DART array. Polarimetric calibration of DART is applied by the channel-based systematic calibration. With the external and internal systematic calibration, the amplitude and phase accuracy of 0.2 dB and between any pair of receivers is guaranteed. With the low residual calibration errors, the polarimetric measurement error of DART is lower than 5% yan2023super.
S2 FAST Observation and Flux Estimation
We conducted a joint observation of DART and FAST at MJD 60559, with the integration time at 1 hour. The FAST frequency band and sampling time are 1-1.5 GHz and 196 us. The 19-beam receiver of FAST adopts orthogonal linear polarisation feeds, the diode noise signal will be injected as the polarization calibrated reference before the target signal recordingli2018fast; jiang2019commissioning; jiang2020fundamental. We conduct the polarimetric calibration by using the PSRCHIVE software.
We estimated the flux density of the pulse based on its SNR, using the radiometer equation:
where K and K/Jy represent the system temperature and gain of FAST jiang2020fundamental, respectively. are the bandwidth and temporal resolution of the observation.
S3 Parkes UWL Observation
We conducted a joint observation of DART and Parkes at MJD 60425. We used Parkes to observe toward the initial position (RA = 18:32:45.0, Dec = -09:12:00.0) for 180 minutes starting from MJD 60425.74625, during which the transient is on the radio-active epoch. We used the standard Parkes backend, Medusa, which applies automatic gain correction for each sub-integration (in a timescale of 0.524288 s). We restored the original result using the scale and offset value recorded in PSRFITS. We extract Parkes UWL subband 0 to 2 (frequency 704 – 1088 MHz) and averaged the entire 384 MHz. We used a time resolution of 0.524288 s, and applied Savitzky-Golay smooth with second-order polynomial fit and a window length of 51 time-samples.
We confirmed that no pulses with an SNR greater than 3 were detected, whereas DART was able to detect them. We contend that the failure to detect these pulses is due to the automatic baseline correction employed by the backend of Parkes.
S4 GTC Observation
We obtained the r-band image of the DART source using the broad-band imaging mode of the OSIRIS on the Gran Telescopio CANARIAS (GTC) on 30 July 2024. The image was co-added using 24 single exposures of 150 s each to avoid sky saturation. A four-point dithering was used to avoid cosmic ray contamination. The total on-source time was 3600 s, and the observation duration lasted about 1.25 hr, including overheads. The observations were conducted under good weather conditions (seeing). The data were first bias-subtracted, flat-fielded, cleaned of cosmic rays, and combined using the ccdproc craig2017astropy and DrizzlePachoffmann2021new packages following standard techniques. The sky background of the entire image was then modeled and subtracted, primarily using the photutilsbradley2016photutils package, following the instructions for modeling the complex 2D background in the Cross-Instrument section of the JWST Data Analysis Tool Notebooks111https://spacetelescope.github.io/jdat_notebooks/notebooks/cross_instrument/background_estimation_imaging/Imaging_Sky_Background_Estimation.html. The detection limit of the co-added image is 24.8 mag, while the upper limit of the background is 26.6 mag.
We searched for sources within a 5′′ radius around the DART coordinates by performing aperture photometry on the co-added image using a 2′′ radius. The top panel in Fig 4 displays the four sources closest to the DART coordinates, along with the reference star. The light curves of these five sources are shown in the bottom, where mag represents the difference between the photometry of each individual exposure and its mean magnitude during the observations. Details of the five targets are provided in Table S1.
| Source | Separation | Magnitude |
| (arcsec) | (mag) | |
| s1 | 1.5 | 21.9 |
| s2 | 4.5 | 24.2 |
| s3 | 5.5 | 24.6 |
| s4 | 5.7 | 25.6 |
| Reference | 8.1 | 20.1 |
Our analysis indicates that no optical counterparts are likely associated with DART J1832-0911. Source shows no significant variability during the observation period. Cross-matching sources through with archival data reveals that only is detected in Pan-STARRS DR2 magnier2020pan, exhibiting a comparable -band magnitude. This suggests that is a typical main-sequence star. Sources to exhibit variability in their light curves. However, is too faint to be reliably detected in individual exposures, and its apparent fluctuations are likely due to detection uncertainties. Moreover, lies outside the plausible location range of the radio transient, further weakening its candidacy. The variable light curve of is characteristic of pulsating variable stars, which are known to produce thermal rather than non-thermal radio emissions christy1966pulsation. As a result, none of the optical sources examined appear to be viable counterparts to DART J1832-0911.
The optical non-detection at the DART coordinates sets an upper limit on its optical counterparts. The radio emissions might originate from a faint dwarf star below our detection threshold, or from other objects not expected to be visible in optical observations, such as a neutron star. In the case of a faint white dwarf, assuming an upper limit of 26.6 mag and a white dwarf radius of , the corresponding effective temperature would be or lower.
The GTC observations place strong constraints on the presence of a luminous dwarf companion of DART J1832-0911 at the estimated distance of 4.5 kpc (with uncertainty of 1.5 kpc). However, we cannot completely rule out the existence of a very faint or cool dwarf below the detection threshold.
S5 Timing Analysis
We collected more than 200 single pulses in the 150–450 MHz range, among which 28 pulses at 425 MHz are with high signal-to-noise ratios. We derive the period through standard pulsar timing methods. Before the timing procedure, we had transferred the local time (in MJD) to the barycenter frame of the solar system. We first use a prior period of 2656.2 s to adopt the single pulses from a continuous time series. Then, we select several pulses with high SNR and take their mean profile as the standard template, to cross-correlate any other pulses to determine the TOA for each pulse. Specifically, we calculated a series of correlations between any single pulse and the standard profile by shifting the bins (i.e. the TOAs) of single pulses. Then we fitted the relation of correlations versus TOAs by a Gaussian function and determined the proper TOA based on fitting result. The TOA errors are derived at a level of Gaussian fitting rather than the conventional , reflecting the fundamental limitations of our timing approach: the time series are extracted from dirty image cubes with a 4 MHz bandwidth and no intra-channel dedispersion, and the finest time resolution is 100 ms against pulse widths of order 100 s. Under these conditions, the threshold provides a more conservative and robust estimate of the TOA uncertainty than the standard convention used in high-precision pulsar timing. Given the large TOA uncertainties inherent to the imaging-based time series and the absence of intra-channel dedispersion, a reliable DM fit from multi-frequency TOAs is not feasible with the DART data alone. We therefore rely on the FAST observation for a precise DM determination, as described in Section S7.
We collect a group of TOAs covering more than two months, a long nulling state prevents us from recording more data. We use the least-square timing method to fit TOAs and derive a period at s. The TOA fitting result is shown in Fig.S3. The period derivative we estimate is s/s. It is important to mention that the fitted TOAs belong to pulses at a central frequency of 425 MHz, because we collected the largest number of pulses at this frequency. Fig.2 shows the P-Pdot diagram of DART J1832-0911. We also present another three reported LPTs in the figure.
S6 Location Constrains and Surrounding Compact Objects
We employ a super-resolution technique to achieve a beam size of 0.5 arcsecond, applying zero-padding interpolation to the sparsely sampled -plane of DART observations. To mitigate systematic errors and phase biases induced by atmospheric turbulence, we align the DART image coordinates with the VLA reference frame using observation from the VLA THOR survey at 1.06 GHz (bihr2016continuum). We select three point sources that can be distinguished from the Galactic radio background in both DART and VLA observations, as marked by blue squares in Fig. S4. The DART coordinate frame is calibrated by fitting these reference points to the VLA frame using the “fit_wcs_from_points” function from the astropy package. The resulting super-resolution DART images, presented in panel (b) of Fig. S4, were obtained at 414 MHz with an integration time of 1,000 s, capturing the LPT emission shown in panel (c).
For precise localization of the target source, we perform a two-dimensional Gaussian fit to the emission region, defined by a 10% (or -0.1 dB) decrease from the peak intensity of the main beam. The fitting results, illustrated in panel (c) of Fig. S4, include contours representing the 68.3%, 95.5%, and 99.7% confidence levels, progressing from innermost to outermost regions. The final estimated position of the target source is RA = 18h32m48.5s, Dec = , with an uncertainty of at the 99.7% confidence level.
Taking the age of SNR G22.7-0.2, we can estimate the escaping velocity of DART J1832-0911 if it was born in the center of this SNR. The 2D transfer speed of a pulsar can be estimated by lyne1994high, where , and are the Right Ascension and Declination in degrees. According to the age limits of the SNR, the 2D transfer velocity of DART J1832-0911 is 176 km/s for 30,000 years and 26 km/s for years. This proper motion speed is quite acceptable in the range of the pulsar proper motion statistics lyne1994high; hobbs2005statistical
DART J1832-0911 lies in the Galactic plane, a region rich with supernova remnants (SNRs) and compact objects. Within our field of view at 425 MHz, there are 7 identified SNRs 222http://snrcat.physics.umanitoba.ca and 39 known compact objects 333https://www.atnf.csiro.au/research/pulsar/psrcat/, including pulsars and high-energy sources. We investigate the character age and DM distance of these compact objects, as indicated in the Fig.5. Objects with DM distances below 2 kpc or above 8 kpc, or characteristic ages exceeding years, are classified as association-unlikely candidates. These thresholds account for a 100% uncertainty in DM distance and a 1000% difference between the SNR’s evolutionary age and the pulsar’s characteristic age, following previous studies gaensler2000large.
Spatial distribution analysis reveals that, within a region centered on DART J1832-0911, the number density of compact objects and SNR is 2.5 deg-2 and 0.4 deg-2, respectively. After excluding unlikely candidates and confirmed associations, the density of potentially associated compact objects drops to 1 deg-2. These densities are significantly lower than in regions with possible coincidental SNR-pulsar overlaps, where compact objects are often suspected to be wandering stars from nearby SNRs. For instance, the PSR J1844-0346 region hosts 10 identified compact objects and at least 13 SNR candidates within approximately 1 deg-2 amenomori2022measurement; kato2024result. Considering the density of the potential associations and the separation distance of the DART source to the SNR center, the probability of an unrelated pulsar coincidentally wandering into the LPT region can be approximated by the fractional area , where is the projected localization radius of the transient (approximately 3 arcminutes). Based on this approximation, the chance of a random pulsar coinciding with the SNR location of DART J1832-0911 is 0.7%.
[…]
Conclusively, we favor the DART J1832-0911 as the potential stellar remains of SNR G22.7-0.2 than any other compact objects.
S7 Dispersion Measure (DM) and Rotation measure (RM)
We measured the DM from DART observation by calculating the time delay between time series at two frequencies, 190 MHz and 205 MHz. The time series spans nearly 5 hours and includes 7 pulses, allowing us to establish a strong correlation between the two series. Nevertheless, the coarse time resolution (500 ms) and large pulse width prevent us from accurately measuring the delay. is determined by fitting the cross-correlation relations between pulse series at two frequencies. For each correlation coefficient, it was derived by shifting the low-frequency series forward step by step.
Using the dispersion relation DM, one can derive the DM value through the , at which the two series are most correlated. Our estimation shows that the time delay between 205 MHz and 190 MHz could be as long as ten seconds, and the corresponding DM is 480136 pc cm-3. The left panel of Fig.S5 shows the correlation coefficients versus the delay time (or DM) and our fitting results.
We measured the DM from FAST spectral by fitting the SNR of de-dispersed pulse at different DMs. The right panel of Fig.S5 shows the SNR of pulse de-dispersed under different DMs.
We measure the RM based on the spectrum observed by FAST. By using the rmfit tool of PSRCHIVE hotan2004psrchive; van2012pulsar, we got the RM estimation of .
The DART data do not allow a robust, independent determination of the RM, because in the DART imaging mode each polarization measurement is integrated over a single 4 MHz-wide band without further intra-channel frequency resolution (Section S1). Therefore, the frequency sampling is insufficient for a reliable fit of polarization position angle as a function of . Moreover, once Stokes and have been averaged over the full 4 MHz band, the information required for channelized derotation is no longer available, so the effect of in-band Faraday rotation cannot be corrected in post-processing within the DART data themselves.
Nevertheless, the RM measured from the FAST burst provides a useful consistency check. Following the standard complex-polarization formalism for Faraday rotation burn1966depolarization; sokoloff1998depolarization; brentjens2005faraday, we define a convenient band-averaged depolarization factor,
where is the fraction of intrinsic linear polarization that survives after averaging over the full channel bandwidth. Here the RM value is and the expected bandwidth is MHz. For representative DART frequencies, we obtain at 150 MHz, at 190 MHz, at 220 MHz, and at 425 MHz. These values correspond to very strong in-band depolarization at the low-frequency end of the DART band and still non-negligible depolarization near 425 MHz. Thus, while the DART data cannot yield a reliable RM, the observed scarcity of linear polarization is compatible with the substantial bandwidth depolarization expected from the RM measured by FAST.
S8 Emission Characters
DART J1832-0911 displays multiple emission characteristics, which we identify as three distinct feature modes: wide-pulse, narrow-pulse, and nulling (or weak emission below DART’s detection threshold) or dwarf-pulse modes. Our observations show no evidence of regular evolution between these features.
In the wide-pulse mode, pulses are typically strong, with widths around 200–250 s, whereas in the narrow-pulse mode, the pulses are much weaker, with widths of roughly 40–100 s, as shown in the pulse stack figures Fig.1 and Fig.S1. For most cases, the pulses remain consistent in both profile and strength, as shown in Fig.S1. However, we observed a significant disruption of this stability on MJD 60439, where flux and profiles dramatically changed, as shown in Fig.S6. This sequence captures a transition from narrow-pulse to wide-pulse mode and back, with two strong wide pulses occurring within the narrow-pulse duration. Predictably, clear circular polarization components appeared at a fixed phase of the corresponding wide-mode pulse. The temporal evolution of the emission—from strong, wide pulses to weak, narrow ones, and eventually to quiescence—suggests a dynamic process of particle accumulation and outward flow within the magnetosphere.
The dwarf pulse of DART J1832-0911 (illustrated in Fig.3) appears to be similar to those observed in some pulsars during their nulling states chen2023strong; yan2024dwarf. These sporadic, brief bursts are attributed to the emission from one or a few “raindrops” of particles generated through pair production in a fragile spark gap within the pulsar’s magnetosphere, close to the neutron star’s surface chen2023strong; yan2024dwarf. That is, a gap model is the underlying mechanism responsible for the radio emission characteristics of DART J1832-0911. Given the extended duration of the pulse components in the LPT, it is possible that such dwarf pulses may be prevalent throughout the entire radiation phase.
S9 Polarization Characteristics
During the emission-active epoch observed from DART, we found poor polarization in most pulses. Our recordings show highly variable polarization, which includes: 1, Absence of Polarization: Both linear and circular polarization were largely undetectable throughout most of the emission cycle. 2, Transient Linear Polarization: A linearly polarized signal occasionally appeared as a short burst. Fig.S7 shows two instances of these bursts, where linear polarization coincided with an intensity spike in the total signal. The majority of pulses, however, show no detectable linear polarization. 3, Phase-Locked Circular Polarization (CP): A stable, circularly polarized component appeared exclusively in the wide-pulse emission mode. Bottom panel of Fig.1 displays the phase-locked CP components recorded. Shaded areas in Fig.S7 and Fig.S6 highlight the CP profile ranges. These pulses include cases with reversed CP signs, cases with no sign change, and cases with no CP emission. These
The scarcity of detectable linear polarization in the DART data is more likely attributable to observational limitations and propagation effects than to intrinsically unpolarized emission. At the low-frequency end of the DART band, the expected RM smearing within a single 4 MHz channel is severe enough to suppress most of the intrinsic linear polarization after frequency averaging. Even near 425 MHz, where the surviving linear-polarization fraction is higher ( for the FAST RM), the remaining signal can still fall below the detection threshold once limited sensitivity is taken into account. The typical noise level of the DART polarization time series is mJy; therefore, for narrow-mode pulses with peak flux densities of – mJy, an intrinsic linear-polarization fraction of would correspond to only – mJy in polarized flux density, well below the noise floor. In addition, the lack of intra-channel dedispersion broadens intrinsically short polarized components and further reduces their apparent peak signal-to-noise ratio. We therefore conclude that the rarity of linear-polarization detections in DART can be naturally explained by the combined effects of bandwidth depolarization, limited sensitivity, and temporal smearing. RM smearing is therefore likely the dominant limitation at the low-frequency end of the DART band, whereas at 425 MHz it remains an important, but not exclusive, contributor.
During the emission-quiescent epoch observed with DART, we detected a bright burst due to the high sensitivity of FAST, with an estimated strength of 94 mJy , which is below DART’s detection threshold. As shown in Fig.3, this pulse was highly linearly polarized, with linear and circular polarization degrees of 94% and 3%, respectively, and a flat position angle (PA) across the burst duration.
Our analysis supports an intrinsic properties of the phase-locked CP, suggesting that an intrinsic spin-down period rather than any extrinsic modulation consequence, e.g. an eclipse influence in a binary system and polarization conversion caused by SNR. During our observation period, both CP-active and CP-quiescent states persisted for several days, as did the wide- and narrow-pulse modes. Known binary systems are unlikely to remain in eclipse for multiple days. Furthermore, as shown in Fig.S6, CP appeared with burst signals even during the prolonged polarization-quiescent period, implying no strict periodicity in the CP-active phases. The phase-locked CP may come from an open-field line region within magnetosphere of a pulsar. The sign change of CP is attributed to the magnetosphere emissions since the line of sight sweeps across the minimum angle between the line of sight and the magnetic axis.
The high LP burst emitted during the DART’s non-detection period demonstrates a strong magnetic field of DART J1832-0911, such near 100% LP emission usually happens in some extreme emission phenomenons, like fast radio bursts (FRBs) ( e.g. michilli2018extreme; andersen2019chime; hilmarsson2021polarization), and some of the giant pulse from radio pulsars jessner2010giant; hankins2016crab. Extremely high linear polarization places severe constraints on the radiation mechanism, only strongly magnetized sources can support relativistic particle bunches that emit highly linearly polarized signals. In addition, a flat PA, similar to some FRBsmichilli2018extreme, indicates an intrinsic limited duration of the emitting process, rather than being influenced by such geometric effects, such as the alignment between the rotation axis and the magnetic axis. One possible emission scenario is that plasma with insufficient density produces nearly 100% linearly polarized emissions with a frozen PA, propagating from the strongly magnetized stellar surface to the exit point, where the wave undergoes minimal rotation. This process occurs on a short-lived timescale lu2019fast.
S10 Model Constraints
In addition to the directly observed proofs mentioned above, we discuss derived constraints about the nature of DART J1832-0911 in this section. According to the radio luminosity and emission duration, all reported LPTs are driven by coherent emission, same as transients like radio pulsars and FRBs, as shown in Fig.S8. However, emission from WD binary system ILT J1101+5521 is predicted also a coherent emitter de2025sporadic. Therefore more constraints should be discussed.
Neutron Star ? We have demonstrated that DART J1832-0911 is likely a young neutron star that has been spun down due to interaction with an SNR fallback disk. Here, we further discuss this source under the assumption that it may be a spin-down powered neutron star. In a spin-down powered scenario, the magnetic field strength can be estimated using G, and the spin-down luminosity is erg s-1 where is the moment of inertia ostriker1969nature; contopoulos2006revised; contopoulos2014new. Taking into account the estimated age of DART J1832-0911, which is constrained by the evolution timescale of the associated SNR to be between and years, we can place limits on the magnetic field strength and spin-down luminosity. Specifically, the magnetic field strength limit is G and the loss rate of rotational energy is erg s-1.
We can compare the theoretical spin-down power to the emission power from the relation erg/s, where is the angular radius of the emitting cone and is the pulse duty cycleszary2014radio. We assume a peak flux of the emission at Jy and the distance at kpc, taking a typical valueszary2014radio of , then the radio luminosity then can be estimated as erg/s.
[…]
where is the moment of inertia of the NS, and are the angles between the spin axis of the NS and rotation axis of the disk, and the magnetic inclination angle of the NS, respectively. The spin evolution of neutron stars assisted by fallback disks is mainly dominated by three torques: the accretion torque , the torques caused by magnetic field-disk interaction and by magnetic dipole radiation . Among these, plays a critical role in the formation of ultra-long-period pulsars.
Moreover, it should be noted that radio emission can only be detected when neutron stars are in the radio ejector phase, which occurs when the inner radius of the disk exceeds the light cylinder radius or the disk becomes passive. Similar to accretion disks in dwarf novae and transient X-ray binaries, the fallback disks would be subject to the thermal ionization instability once the outermost regions of the disk cool sufficiently menou2001stability. Under these conditions, the transport of material along the disk becomes inefficient, resulting in a passive disk and the cessation of its activity. The minimum temperature required for a disk to remain active is approximately 300 K ertan2009evolution. Here we evolved one million systems, assuming that the initial distributions of , , and follow as: , , , and . The simulated results of and at yr and yr, corresponding to the age estimate of SNR G22.7-0.2, are shown as green and purple shaded areas in Fig.2. It is clear that these LPTs can be well covered by the simulated distributions of ejectors, which may support the hypothesis that such LPTs may originate from interactions between young neutron stars and fallback discs.
Magnetic White Dwarf ? The white dwarf model faces challenges in explaining DART J1832-0911s properties but cannot be entirely ruled out. The radio-limited brightness temperature provides constraints on the source quantities by examining the relationship between brightness temperatures and the size of the emitting region, applicable to both thermal and coherent (non-thermal) emitters. Using the Rayleigh-Jeans approximation as a general definition of brightness temperature:
where denotes the size of the emitting region (rather than the stellar radius). At MHz and a distance range of kpc (bracketing estimates from the YMW16, NE2001, and NE2025 electron-density models), the resulting as a function of emitting region size is depicted in Fig. S9.
For any emitting region size comparable to a white dwarf radius (or much smaller), the inferred K is characteristic of coherent emission, consistent with the broader LPT population as shown in Fig.S8. Since coherent emission is expected in both neutron star and magnetic white dwarf models marsh2016radio; pelisoli20235; de2025sporadic; hurley20242; dong2025chime; bloot2025strongly, the alone does not discriminate between the two scenarios. However, the required brightness temperature still provides constraints. In confirmed WD-binary LPTs such as ILT J1101+5521, the coherent radio emission arises from a large-scale interaction region between the two stars, with the emitting region on the order of the companion stellar radius () de2025sporadic, yielding K. By contrast, the isolated nature of DART J1832-0911 implies that any WD model would require the coherent emission to originate from a compact region such as the WD magnetic pole (), pushing to K. Such extreme brightness temperature implies a highly ordered and energetic particle beam, requiring a magnetic field strength well in excess of that observed in any known radio-emitting white dwarf ( G for AR Sco marsh2016radio), and more consistent with neutron star magnetospheric conditions.
Another constraint comes from the energy budget and the characteristic cooling timescale. Before the crystallization, the cooling timescale of a hot WD is dominated by losing its thermal energy through black body radiation. Theoretically, the cooling time is a function of the WD luminosity , WD mass , core atomic weight . The approximate relation can be expressed as mestel1952theory; althaus2010evolutionary
where and are the solar mass and luminosity. The WD luminosity is . Assuming a typical WD with pure carbon core, then , the mass and radius are kepler2007white and km provencal1998testing, respectively. We then obtained a cooling time should be longer than years as K.
Since WDs could be also spin-down powered dominated marsh2016radio, we can estimate the characteristic age of a white dwarf candidate using the spin-down mechanism, which is given by:
For such a radio-emitting WD , similar to the neutron star case geroyannis2000spin; buckley2017polarimetric, we calculate the spin-down luminosity by adjusting for the WD’s larger moment of inertia, typically g for a WDgeroyannis2000spin. Assuming that the radio emission of a WD is beamed like that of a pulsar, we then take the emission power erg/s as its rotational energy-loss rate. Then we derive the period derivative is s/s. This gives a characteristic age of years, which is consistent with the lower limit cooling time scale based on the optical observation. However, such an analysis is parameter-dependent and serves only as a description of typical cases.
In conclusion, if a spin-down powered white dwarf is responsible for DART J1832-0911, it should be an extremely strongly magnetized, old and cold one.
Declarations
Conflict of interest. The authors declare that they have no conflict of interest.
Author contributions. D. Li, M. Yuan, and L. Wu conceived and designed the study. J.-Y. Yan developed the detection pipeline/method and carried out the search. X.-L. Lv coordinated the observations and collected the data. All the other authors contributed to calibration and validation of the data, interpretation of the results, and preparation of the manuscript.
Data availability. The data that support the findings of this study are available from the corresponding author upon reasonable request.